OpenStudy (anonymous):

lim 1-sqrtx/1-x as x goes to 1

myininaya (myininaya):

factor 1-x :)

myininaya (myininaya):

or

myininaya (myininaya):

IF you don't like that rationalize the numerator

myininaya (myininaya):

I prefer the first way :)

myininaya (myininaya):

\[a-b=(\sqrt{a}-\sqrt{b})(\sqrt{a}+\sqrt{b})\]

OpenStudy (anonymous):

could explain it to me I think I got stuck after that part.

myininaya (myininaya):

Did you factor 1-x? What did you get?

OpenStudy (anonymous):

is it 1/x I'm not sure

myininaya (myininaya):

Ok since you can't factor 1-x using the formula I gave Then do you know how to rationalize the numerator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(1-\sqrt{x}) / (1-\sqrt{x}) (1+\sqrt{x})\]

OpenStudy (anonymous):

Now solve..:)

myininaya (myininaya):

like for example pretend I wanted to factor 9-x using the formula I gave it would be \[(3-\sqrt{x})(3+\sqrt{x})\]

myininaya (myininaya):

But you can always go the rationalizing the numerator way for this one Just multiply both top and bottom by top's conjugate

OpenStudy (anonymous):

so on top would be \[(1+\sqrt{x})(\sqrt{x-1})\] right?

myininaya (myininaya):

No your second factor is a bit messed up

OpenStudy (anonymous):

or would it be x-1

myininaya (myininaya):

\[4-x=(2-\sqrt{x})(2+\sqrt{x})\] \[9-x=(3-\sqrt{x})(3+\sqrt{x})\] \[16-x=(4-\sqrt{x})(4+\sqrt{x})\] \[25-x=(5-\sqrt{x})(5+\sqrt{x})\] Do you see I'm just using the formula above: \[a-b=(\sqrt{a}-\sqrt{b})(\sqrt{a}+\sqrt{b})\]

myininaya (myininaya):

So \[1-x=(1-\sqrt{x})(1+\sqrt{x})\]

myininaya (myininaya):

Got it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so that would be on top and bottom Right?

myininaya (myininaya):

ok where you have 1-x you will have \[(1-\sqrt{x})(1+\sqrt{x})\]

myininaya (myininaya):

which is on bottom right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

Ok so does anything cancel?

OpenStudy (anonymous):

yes the \[1-\sqrt{x}\] which 1 is on the bottom and top right?

myininaya (myininaya):

\[\lim_{x \rightarrow 1}\frac{1-\sqrt{x}}{1-x}=\lim_{x \rightarrow 1}\frac{1-\sqrt{x}}{(1-\sqrt{x})(1+\sqrt{x})}\] You have the factor \[1-\sqrt{x}\] on both top and bottom so cancel it

OpenStudy (anonymous):

ok so all I will have left is 1+sqrt x on bottom which is 1+the sqrt of 1 =2 Right? so the anserw is 1/2

myininaya (myininaya):

that's right

OpenStudy (anonymous):

thanks myininaya